# On the Convergence Rates of Federated Q-Learning across Heterogeneous Environments

### v1

@@ 5 Main Results
$\bar{Q}_{t+1}=\frac{1}{K}\sum_{k=1}^{K}\left((1-\lambda)Q_{t}^{k}+\lambda(R+\gamma\widetilde{P}^{k}_{t}V_{t}^{k})\right),$
@@
To show the convergence of $\left\|{\Delta_{t+1}}\right\|_{\infty}$, we bound each of the three terms in the right-hand-side of equation (1). The following lemma is a coarse upper bound of errors.
@@
With the choice of $Q_{0}$ in Lemma 2, the first term in equation (1) can be bounded as $\left\|{(1-\lambda)^{t+1}\Delta_{0}}\right\|_{\infty}\leq(1-\lambda)^{t+1}\frac{1}{1-\gamma}$. In addition, as detailed in the proof of Lemma 4 and Theorem 1, the boundedness in Lemma 2 enables us to bound the second term in equation (1) via invoking the Hoeffding’s inequality. It remains to bound the third term in equation (1), for which we follow the analysis roadmap of Woo et al. (2023) by a two-step procedure that is described in Lemma 3 and Lemma 4. Let
@@
Both Lemma 3 and Lemma 4 are non-trivial adaptations of the characterization in the analysis of Woo et al. (2023) due to lack of common optimal action for any given state when environments are heterogeneous.
To bound the $\ell_{\infty}$ norm of the third term in equation (1), we first invoke Lemma 3, followed by Lemma 4. It is worth noting that directly applying Lemma 4 can also lead to a valid error bound yet the resulting bound will not decay as $T$ increases with proper choice of stepsizes.
@@
Choose $E-1\leq\frac{1-\gamma}{4\gamma\lambda}$ and $\lambda\leq\frac{1}{E}$. For any $\delta\in(0,\frac{1}{3})$, with probability at least $1-3\delta$, it holds that
$\left\|{\Delta_{T}}\right\|_{\infty}$ $\leq\frac{4}{(1-\gamma)^{2}}\exp\left\{-\frac{1}{2}\sqrt{(1-\gamma)\lambda T}\right\}+\frac{2\gamma^{2}}{(1-\gamma)^{2}}(6\lambda^{2}(E-1)^{2}+\lambda(E-1))\kappa$
$\quad+\left(\frac{12\gamma^{2}\lambda}{(1-\gamma)^{2}}\sqrt{E-1}+\frac{2\gamma^{2}\sqrt{\lambda}}{(1-\gamma)^{2}}\right)\sqrt{\lambda(E-1)\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}$
$\quad+\frac{2\gamma}{(1-\gamma)^{2}}\sqrt{\frac{1}{K}\lambda\log\frac{|S||A|TK}{\delta}}.$
The first term of Theorem 1 is the standard error bound in the absence of environmental heterogeneity and sampling noises. The second term arises from environmental heterogeneity. It is clear that when $E=1$, the environmental heterogeneity does not negatively impact the convergence. The last two terms result from the randomness in sampling.
@@
It is common to choose the stepsize $\lambda$ based on the time horizon $T$. Let $\lambda=g(T)$ be a non-increasing function of $T$. As long as $\lambda=g(T)$ decay in $T$, terms 2-4 in Theorem 1 will go to 0 as $T$ increases. In addition, when $\lambda=\omega(1/T)$, the first term will decay to 0.
There is a tradeoff in the convergence rates of the first term and the remaining terms – the slower $\lambda$ decay in $T$ leads to faster decay in the first term but slower in the remaining terms. Forcing these terms to decay around the same speed lead to slow overall convergence. Corollary 1 follows immediately from Theorem 1 via carefully choosing $\lambda$ to balance the decay rates of different terms.
@@
Choose $(E-1)\leq\min\frac{1}{\lambda}\{\frac{\gamma}{1-\gamma},\frac{1}{K}\}$, and $\lambda=\frac{4\log^{2}(TK)}{T(1-\gamma)}$. Let $T\geq E$. For any $\delta\in(0,\frac{1}{3})$, with probability at least $1-3\delta$,
$\left\|{\Delta_{T}}\right\|_{\infty}$ $\leq\frac{4}{(1-\gamma)^{2}TK}+\frac{36}{(1-\gamma)^{3}}\frac{\log(TK)}{\sqrt{TK}}\sqrt{\log\frac{|S||A|TK}{\delta}}+\frac{56\log^{2}(TK)}{(1-\gamma)^{3}}\frac{E-1}{T}\kappa.$
Remark 2 (Partial linear speedup and the negative impacts of $E>1$).
Intuitively, both terms 1 and 2 decay as if there are $TK$ iterations – a linear speedup. In fact, the decay rate of the sampling noises in Corollary 1, with respect to $TK$, is the minimax optimal up to polylog factors (Vershynin, 2018). The decay of the third term is controlled by environmental heterogeneity when $E>1$. In sharp contrast to the homogeneous settings, larger $E$ significantly slows down the convergence of this term.
We show in the next subsection that this slow convergence is fundamental.
5.2 On the fundamentals of Convergence Slowing Down for $E>1$.
@@
Let $Q_{0}=\bm{0}$. For any even $K\geq 2$, there exist a collection of $\{({\mathcal{S}},{\mathcal{A}},{\mathcal{P}}^{k},R,\gamma):~{}k\in[K]\}$ such that, for any synchronization period $E$ and time-invariant stepsize $\lambda\leq\frac{1}{1+\gamma}$,
$\left\|{\Delta_{T}}\right\|_{\infty}\geq\Omega\left(E/T\right),$
when $T/E\in{\mathbb{N}}$ and $T\geq E\cdot\exp\left\{-W_{-1}\left(-\frac{1-\gamma}{2(1+\gamma)}\right)\right\},\text{ where }W_{-1}\text{ is the Lambert $W$ function}$.
Proof Sketch. Below we discuss the key ideas and provide the proof sketch of Theorem 2.
@@
Following Algorithm 1, we obtain the following recursion between two synchronization rounds:
$\Delta_{(r+1)E}=\bar{A}^{(E)}\Delta_{rE}+\left(\left(I-\bar{A}^{(E)}\right)-\left(I+\bar{A}^{(1)}+\dots\bar{A}^{(E-1)}\right)\left(I-\bar{A}^{(1)}\right)\right)Q^{*},$ (9)
where $\bar{A}^{(\ell)}\triangleq\frac{1}{K}\sum_{k=1}^{K}(A^{k})^{\ell}$ and $A^{k}\triangleq(1-\lambda)I+\lambda\gamma P^{k}$. While the first term on the right-hand side of equation (9) decays rapidly to zero, the second term is non-vanishing due to environment heterogeneity for $E\geq 2$. Specifically, to ensure the rapid decay of the first term, it is necessary to select a stepsize $\lambda=\widetilde{\Omega}(\frac{1}{rE})$. However, this choice results in the dominating residual error from the second term, which increases linearly with $\lambda E=\widetilde{\Omega}(\frac{1}{r})$.
Next, we instantiate the analyses by constructing the set ${\mathcal{P}}^{k}$ over a pair of states and an even number of clients with
@@
where $\bar{P}=\frac{1}{2}\mathbf{11}^{\top}$, $\alpha_{\ell}\triangleq\frac{1}{2}(\nu_{1}^{\ell}+\nu_{2}^{\ell})$, $\beta_{\ell}\triangleq\nu_{2}^{\ell}$, $\nu_{1}\triangleq 1-(1+\gamma)\lambda$, and $\nu_{2}\triangleq 1-(1-\gamma)\lambda$. For this instance of ${\mathcal{P}}_{k}$, the error evolution  equation (9) reduces to $\Delta_{(r+1)E}=\left(\alpha_{E}(I-\bar{P})+\beta_{E}\bar{P}\right)\Delta_{rE}+\kappa_{E}(I-\bar{P})Q^{*}$ with $\kappa_{E}\triangleq-\frac{\gamma}{2}\left(\frac{1-\nu_{2}^{E}}{1-\gamma}-\frac{1-\nu_{1}^{E}}{1+\gamma}\right)$, which further yields the following full error recursion:
$\Delta_{rE}=\left(\alpha_{E}^{r}(I-\bar{P})+\beta_{E}^{r}\bar{P}\right)\Delta_{0}+\frac{1-\alpha_{E}^{r}}{1-\alpha_{E}}\kappa_{E}(I-\bar{P})Q^{*}.$
Starting from $Q_{0}=0$, the error can be decomposed into
$\Delta_{rE}=\beta_{E}^{r}\bar{P}Q^{*}+\left(\alpha_{E}^{r}+\frac{1-\alpha_{E}^{r}}{1-\alpha_{E}}\kappa_{E}\right)(I-\bar{P})Q^{*}.$ (11)
The two terms of the error are orthogonal and both non-vanishing. Therefore, it remains to lower bound the maximum magnitude of two coefficients irrespective of the stepsize $\lambda$. To this end, we analyze two regimes of $\lambda$ separated by a threshold $\lambda_{0}\triangleq\frac{\log r}{(1-\gamma)rE}$:
  • Slow rate due to small stepsize when $\lambda\leq\lambda_{0}$. Since $\beta_{E}^{r}$ decreases as $\lambda$ increases,
    $\beta_{E}^{r}\geq(1-(1-\gamma)\lambda_{0})^{rE}=\left(1-\frac{\log r}{rE}\right)^{rE}\gtrsim\frac{1}{r}.$
@@
    $\left|\frac{\kappa_{E}}{1-\alpha_{E}}\right|\geq\gamma^{2}\frac{\lambda(E-1)}{4}\gtrsim\gamma^{2}\frac{\log r}{(1-\gamma)r},\qquad\left(1+\left|\frac{\kappa_{E}}{1-\alpha_{E}}\right|\right)\alpha_{E}^{r}\leq\frac{1}{(1-\gamma^{2})r}.$
We conclude that at least one component of the error in equation (11) must be slower than the rate $\Omega(1/r)$.
@@
The explicit calculations are based on a set ${\mathcal{P}}^{k}$ over a pair of states. Nevertheless, the evolution equation (9) is generally applicable. Similar analyses can be extended to scenarios involving more than two states, provided that the sequence of matrices $\bar{A}^{(\ell)}$ is simultaneously diagonalizable. For instance, the construction of the transition kernels in equation (10) can be readily extended to multiple states if the set ${\mathcal{S}}$ can be partitioned into two different classes. The key insight is the non-vanishing residual on the right-hand side of equation (9) when $E\geq 2$ due to the environment heterogeneity.

### v2

@@ 5 Main Results
$\bar{Q}_{t+1}=\frac{1}{K}\sum_{k=1}^{K}\left((1-\lambda)Q_{t}^{k}+\lambda(R+\gamma\widetilde{P}^{k}_{t}V_{t}^{k})\right),$ (4)
@@
To show the convergence of $\left\|{\Delta_{t+1}}\right\|_{\infty}$, we bound each of the three terms in the right-hand side of (1). The following lemma is a coarse error upper bound.
@@
With the choice of $Q_{0}$ in Lemma 2, the first term in (1) can be bounded as $\left\|{(1-\lambda)^{t+1}\Delta_{0}}\right\|_{\infty}\leq(1-\lambda)^{t+1}\frac{1}{1-\gamma}$. In addition, as detailed in the proof of Lemma 4 and Theorem 1, the boundedness in Lemma 2 enables us to bound the second term in (1) via invoking the Hoeffding’s inequality. It remains to bound the third term in (1), for which we follow the analysis roadmap of Woo et al. (2023) by a two-step procedure that is described in Lemma 3 and Lemma 4. Let
@@
To bound the $\ell_{\infty}$ norm of the third term in (1), we first invoke Lemma 3, followed by Lemma 4. It is worth noting that directly applying Lemma 4 can also lead to a valid error bound yet the resulting bound will not decay as $T$ increases with proper choice of stepsizes.
Both Lemma 3 and Lemma 4 are non-trivial adaptations of the approach in Woo et al. (2023) due to the absence of a common optimal action for any given state in heterogeneous environments. Moreover, in the homogeneous setting, each agent draws samples from the same true transition distribution, allowing concentration inequalities to bound the discrepancy between the true distribution and sampled estimates. However, this line of reasoning does not go through in the presence of environmental heterogeneity. When $\kappa>0$, each of the $K$ agents has its own transition distribution, and the discrepancy is captured by the environmental heterogeneity parameter $\kappa$.
@@
Choose $E-1\leq\frac{1}{\lambda}\min\{\frac{1-\gamma}{4\gamma},\frac{1}{K}\}$ and $\lambda\leq\frac{1}{E}$. For any $\delta\in(0,\frac{1}{3})$, with probability at least $1-3\delta$, it holds that
$\left\|{\Delta_{T}}\right\|_{\infty}$ $\leq\frac{4}{(1-\gamma)^{2}}\exp\left\{-\frac{1}{2}\sqrt{(1-\gamma)\lambda T}\right\}+\frac{14\gamma^{2}}{(1-\gamma)^{2}}\lambda(E-1)\kappa+\frac{16}{(1-\gamma)^{2}}\sqrt{\frac{\lambda}{K}\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}.$
The first term of Theorem 1 is the standard error bound in the absence of environmental heterogeneity and sampling noises. The second term arises from environmental heterogeneity. It is clear that when $E=1$, the environmental heterogeneity does not negatively impact the convergence. The last term results from the randomness in sampling.
@@
It is common to choose the stepsize $\lambda$ based on the time horizon $T$. Let $\lambda=g(T)$ be a non-increasing function of $T$, and other parameters be fixed with respect to $T$. As long as $\lambda=g(T)$ decay in $T$, terms 2 and 3 in Theorem 1 will go to 0 as $T$ increases. In addition, when $\lambda=\omega(1/T)$, the first term will decay to 0. Conversely, the convergence bounds in Zhang et al. (2024) and Wang et al. (2023) do not decay to 0.
There is a tradeoff in the convergence rates of the first term and the remaining terms – the slower $\lambda$ decay in $T$ leads to faster decay in the first term but slower in the remaining terms. Forcing these terms to decay around the same speed leads to slow overall convergence. Corollary 1 follows immediately from Theorem 1 via carefully choosing $\lambda$ to balance the decay rates of different terms.
@@
Choose $(E-1)\leq\min\frac{1}{\lambda}\{\frac{1-\gamma}{4\gamma},\frac{1}{K}\}$, and $\lambda=\frac{4\log^{2}(TK)}{T(1-\gamma)}$. Let $T\geq E$. For any $\delta\in(0,\frac{1}{3})$, with probability at least $1-3\delta$,
$\left\|{\Delta_{T}}\right\|_{\infty}$ $\leq\frac{4}{(1-\gamma)^{2}TK}+\frac{32}{(1-\gamma)^{2.5}}\frac{\log(TK)}{\sqrt{TK}}\sqrt{\log\frac{|{\mathcal{S}}||{\mathcal{A}}|TK}{\delta}}+\frac{56\log^{2}(TK)}{(1-\gamma)^{3}}\frac{E-1}{T}\kappa.$
Remark 2.
Intuitively, both terms 1 and 2 decay as if there are $TK$ iterations. In fact, the decay rate of the sampling noises in Corollary 1, with respect to $TK$, is the minimax optimal up to polylog factors (Vershynin, 2018). The decay of the third term is controlled by environmental heterogeneity when $E>1$. In sharp contrast to the homogeneous settings, larger $E$ significantly slows down the convergence of this term. We show in the next subsection that this slow convergence is fundamental.
Remark 3 (Communication cost and convergence).
From Corollary 1, by choosing $E=\widetilde{\Theta}(\sqrt{T})$, and other parameters are fixed with respect to $T$, we can reach the same error bound of $\widetilde{\mathcal{O}}(1/\sqrt{T})$ with communication cost of $\widetilde{{\mathcal{O}}}(\sqrt{T})$, which is better than $\widetilde{{\mathcal{O}}}(T)$.
Corollary 2.
Choose $E-1\leq\frac{1}{\lambda}\min\{\frac{1-\gamma}{4\gamma},\frac{1}{K}\}$ and $\lambda\leq\frac{1}{E}$, and define $x_{1}=\frac{4096\log\frac{|\mathcal{S}||\mathcal{A}|K}{\delta}\log^{2}(\frac{(1-\gamma)^{2}\epsilon}{8})}{K(1-\gamma)^{5}\epsilon^{2}}$$x_{2}=\frac{168\kappa(E-1)\gamma^{2}}{(1-\gamma)^{3}\epsilon}\log^{2}(\frac{(1-\gamma)^{2}\epsilon}{12}),$ and $x_{3}=\frac{9216}{K(1-\gamma)^{5}\epsilon^{2}}\log\frac{|\mathcal{S}||\mathcal{A}|K}{\delta}\log^{2}(\frac{(1-\gamma)^{2}\epsilon}{12})$,
  • When $\kappa=0\text{ or }E=1,$ for any $\delta\in(0,\frac{1}{3})$, with probability at least $1-3\delta$, it holds that
    $\left\|{\Delta_{T}}\right\|_{\infty}\leq\epsilon,$
    when $T\geq\exp\{-W_{-1}(-\frac{1}{x_{1}})\}$, and $\lambda=\frac{\epsilon^{2}(1-\gamma)^{4}K}{2304\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}$, $\text{where }W_{-1}\text{ is the Lambert $W$}$ function. The resulting sample complexity is $\widetilde{\mathcal{O}}\left(\frac{|{\mathcal{S}}||{\mathcal{A}}|}{K(1-\gamma)^{5}\epsilon^{2}}\right)$.
  • When $\kappa>0\text{ and }E>1,$
    • If $K(E-1)\geq\frac{55\log\frac{|{\mathcal{S}}||{\mathcal{A}}|K}{\delta}}{\kappa\gamma^{2}\epsilon(1-\gamma)^{2}}$, for any $\delta\in(0,\frac{1}{3})$, with probability at least $1-3\delta$, it holds that
      $\left\|{\Delta_{T}}\right\|_{\infty}\leq\epsilon,$
      when $T\geq\exp\{-W_{-1}(-\frac{1}{x_{2}})\}$ and $\lambda=\frac{\epsilon(1-\gamma)^{2}}{42\kappa(E-1)\gamma^{2}}$ . The sample complexity is $\widetilde{\mathcal{O}}\left(\frac{\kappa|{\mathcal{S}}||{\mathcal{A}}|E}{(1-\gamma)^{3}\epsilon}\right)$.
    • If $K(E-1)\leq\frac{55\log\frac{|{\mathcal{S}}||{\mathcal{A}}|K}{\delta}}{\kappa\gamma^{2}\epsilon(1-\gamma)^{2}}$, for any $\delta\in(0,\frac{1}{3})$, with probability at least $1-3\delta$, it holds that
      $\left\|{\Delta_{T}}\right\|_{\infty}\leq\epsilon,$
      when $T\geq\exp\{-W_{-1}(-\frac{1}{x_{3}})\}$ and $\lambda=\frac{\epsilon^{2}(1-\gamma)^{4}K}{2304\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}$. The sample complexity is $\widetilde{\mathcal{O}}\left(\frac{|{\mathcal{S}}||{\mathcal{A}}|}{K(1-\gamma)^{5}\epsilon^{2}}\right)$.
Remark 4 (Sample complexity on $K$ and $E$ and conditional linear speedup.).
From Corollary 2, we can conclude that when the setting is homogeneous (i.e., $\kappa=0$) or the agents communicate every step (i.e., $E=1$), the sample complexity $\widetilde{\mathcal{O}}\left(\frac{|{\mathcal{S}}||{\mathcal{A}}|}{K(1-\gamma)^{5}\epsilon^{2}}\right)$ matches the one in Woo et al. (2023). On the other hand, when the setting is heterogeneous (i.e., $\kappa>0$) and $E>1$, it is evident that if the total computation steps per synchronization are sufficiently small, i.e., $K(E-1)\leq\widetilde{\mathcal{O}}((\kappa\epsilon)^{-1}(1-\gamma)^{-2})$, the sample complexity also matches the one in the homogeneous setting, where there is a linear speedup. Otherwise, the sample complexity $\widetilde{\mathcal{O}}\left(\frac{\kappa|{\mathcal{S}}||{\mathcal{A}}|E}{(1-\gamma)^{3}\epsilon}\right)$ increases with $E$, meaning that multiple local rounds only consume more samples (i.e., $E$-times more) samples without achieving linear speedup.
5.2 On the Fundamentals of Convergence Slowdown for $E>1$ in Heterogeneous Environments.
@@
Let $Q_{0}=\bm{0}$. For any even $K\geq 2$, there exists a collection of $\{({\mathcal{S}},{\mathcal{A}},{\mathcal{P}}^{k},R,\gamma):~k\in[K]\}$ where $|{\mathcal{S}}|=2$, $|{\mathcal{A}}|=1$, and $R:=\begin{bmatrix}r_{1}\\ r_{2}\end{bmatrix}$ fixed for all MDPs in the collection, such that, for $E\geq 2$ and time-invariant stepsize $\lambda\leq\frac{1}{1+\gamma}$,
$\left\|{\Delta_{T}}\right\|_{\infty}\geq c_{R}\frac{E}{(1-\gamma)T},$
when $T/E\in{\mathbb{N}}$ and $T\geq E\cdot\max\left\{\exp\{\frac{4E(\gamma+2)}{(1+\gamma)\gamma^{2}(E-1)}\},\exp\left\{-W_{-1}\left(-\frac{1-\gamma}{2(1+\gamma)}\right)\right\}\right\},\text{where }W_{-1}\text{ is the Lambert $W$}$ function, $c_{R}=\frac{1}{2}\min\left\{\frac{\left|r_{1}+r_{2}\right|}{e},\left|r_{1}-r_{2}\right|\right\}$ when $r_{1}\not=r_{2}$ and $c_{R}=\frac{\left|r_{1}+r_{2}\right|}{2e}$ otherwise.
Proof Sketch. Below we provide the proof sketch of Theorem 2. The full proof is deferred to Appendix F.
@@
Following Algorithm 1, we let $z$ denote the $z$-th synchronization round, and obtain the following recursion between two synchronization rounds:
$\Delta_{({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}+1)E}=\bar{A}^{(E)}\Delta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}E}+\left(\left(I-\bar{A}^{(E)}\right)-\left(I+\bar{A}^{(1)}+\dots\bar{A}^{(E-1)}\right)\left(I-\bar{A}^{(1)}\right)\right)Q^{*},$ (10)
where $\bar{A}^{(\ell)}\triangleq\frac{1}{K}\sum_{k=1}^{K}(A^{k})^{\ell}$ and $A^{k}\triangleq(1-\lambda)I+\lambda\gamma P^{k}$. While the first term on the right-hand side of (10) decays rapidly to zero, the second term is non-vanishing due to environment heterogeneity for $E\geq 2$. Specifically, to ensure the rapid decay of the first term, it is necessary to select a stepsize $\lambda=\widetilde{\Omega}(\frac{1}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}E})$. However, this choice results in the dominating residual error from the second term, which increases linearly with $\lambda E=\widetilde{\Omega}(1/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z})$.
Next, we instantiate the analyses by constructing the set ${\mathcal{P}}^{k}$ over two states and an even number of clients with
@@
where $\bar{P}=\frac{1}{2}\mathbf{11}^{\top}$, $\alpha_{\ell}\triangleq\frac{1}{2}(\nu_{1}^{\ell}+\nu_{2}^{\ell})$, $\beta_{\ell}\triangleq\nu_{2}^{\ell}$, $\nu_{1}\triangleq 1-(1+\gamma)\lambda$, and $\nu_{2}\triangleq 1-(1-\gamma)\lambda$. For this instance of ${\mathcal{P}}_{k}$, the error evolution (10) reduces to $\Delta_{({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}+1)E}=\left(\alpha_{E}(I-\bar{P})+\beta_{E}\bar{P}\right)\Delta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}E}+\kappa_{E}(I-\bar{P})Q^{*}$ with $\kappa_{E}\triangleq-\frac{\gamma}{2}\left(\frac{1-\nu_{2}^{E}}{1-\gamma}-\frac{1-\nu_{1}^{E}}{1+\gamma}\right)$, which further yields the following full error recursion:
$\Delta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}E}=\left(\alpha_{E}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}}(I-\bar{P})+\beta_{E}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}}\bar{P}\right)\Delta_{0}+\frac{1-\alpha_{E}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}}}{1-\alpha_{E}}\kappa_{E}(I-\bar{P})Q^{*}.$
Starting from $Q_{0}=\bm{0}$, the error can be decomposed into
$\Delta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}E}=\beta_{E}^{z}\bar{P}Q^{*}+\left(\alpha_{E}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}}+\frac{1-\alpha_{E}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}}}{1-\alpha_{E}}\kappa_{E}\right)(I-\bar{P})Q^{*}.$ (12)
The two terms of the error are orthogonal and both non-vanishing. Therefore, it remains to lower bound the maximum magnitude of two coefficients irrespective of the stepsize $\lambda$.
To this end, we analyze two regimes of $\lambda$ separated by a threshold $\lambda_{0}\triangleq\frac{\log(T/E)}{(1-\gamma)T}$:
  • Slow rate due to small stepsize when $\lambda\leq\lambda_{0}$. Since $\beta_{E}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}}$ decreases as $\lambda$ increases,
    $\beta_{E}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}}\geq(1-(1-\gamma)\lambda_{0})^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}E}=\left(1-\frac{\log{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}E}\right)^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}E}{\geq\frac{E}{eT}}.$
@@
    $\left|{\alpha_{E}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}}+\frac{1-\alpha_{E}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z}}}{1-\alpha_{E}}\kappa_{E}}\right|\geq\frac{E}{(1-\gamma)T}.$
We conclude that at least one component of the error in (12) must be slower than the rate $\Omega(E/T)$.
@@
The explicit calculations are based on a set ${\mathcal{P}}^{k}$ over a pair of states. Nevertheless, the evolution (10) is generally applicable. Similar analyses can be extended to scenarios involving more than two states, provided that the sequence of matrices $\bar{A}^{(\ell)}$ is simultaneously diagonalizable. For instance, the construction of the transition kernels in (11) can be readily extended to multiple states if the set $\mathcal{S}$} can be partitioned into two different classes. The key insight is the non-vanishing residual on the right-hand side of (10) when $E\geq 2$ due to the environment heterogeneity.
5.3 Discussion on Time-varying Stepsize
Although using time-varying stepsize is common and simple when implementing the algorithm, it is not easy to transfer from current time-invariant stepsize analysis to time-varying stepsize analysis. This is because in the time-invariant stepsize analysis we are dealing with a function of one variable, however, in the time-varying case, we are dealing with a function of $T$ variables.
For example, in our lower bound analysis, we picked a threshold $\lambda_{0}$ and showed that no matter $\lambda$ is greater or smaller than $\lambda_{0}$, the convergence rate is greater than $\Theta_{R}\left(E/((1-\gamma)T)\right)$, and we can claim we have covered all the cases. However, for time-varying stepsize, the number of stepsizes is $T$, and it is not easy to generalize a similar result by just considering several cases because each stepsize gives an additional dimension. Even if we know the sequence is decaying, without specifying a particular family of stepsizes, it is not possible to divide it into several cases as we did for time-invariant stepsize.
We conjecture that both approaches lead to comparable residual error levels over extended training. For example, the stepsizes used in Figure 4(a) and Figure 4(b) are $\frac{1}{\sqrt{T}}$ and $\frac{1}{\sqrt{t+1}}$, respectively. While the time-decaying stepsize appears to have faster initial convergence due to its larger values, we observe that as $t$ increases, the convergence rates of the two strategies seem to align, suggesting a similar asymptotic behavior.